OpenStudy (anonymous):

Solve the equation by completing the square. Round to the nearest hundredth if necessary. x^2 + 10x = 18

sam (.sam.):

These are ancient to me but I'll try, Subtract both sides by 18 x^2+10x-18=0 Divide 'b' by 2, simplify x^2 and 5x you get (x+5)^2, square '5', you get 25, \((x+5)^2-25-18=0\)

sam (.sam.):

Finish it up

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

Wait, I forgot to list the answer choices. The choices are: A. -11.56, 1.56 B. 11.56, 1.56 C. -11.56, -1.56 D. 11.56, -1.56 I'll try and finish the equation now: So I would distribute the exponent, so that makes x^2 + 25 - 25 - 18 = 0 then I would do the edition/subtraction: x^2 - 18 = 0 and now I would add 18 to both sides? x^2 = 18 but then x would equal 4.24......so that's wrong